The height of a cylindrical tank, containing 15 lb of gas at 70 psia and 115 Farenheit, IS TRIPLE ITS DIAMETER; the gas constant is R = 45 ft lbf/lbm R. find.

a) the tankdiameter in ft.

b)the tank height in ft.


1
Expert's answer
2021-11-19T10:51:40-0500

W=15 lbm, P= 70 psia, T=115 F, H=3D, R = 45 ft lbf/lbm°R

Volume of cylireler

"V=\frac{π}{4} D^2\times H=\frac{π}{4}D^2\times3D=\frac{3π}{4}D^3\\

V=\frac{3}{4}πD^3"

1F=459.67R

1ft=12inches

So, T=115F=115+459.67=574.67R

"Psia=\frac{lbf}{in^2}=\frac{lbf}{(\frac{1}{12})^2ft^2}"

By ideal gas equation

PV=mrt

"\frac{70}{ (\frac{1}{12})^2}( \frac{lbf}{ft^2})\times \frac {3π}{4}D^3(ft^3)=15(lbm)\times45(\frac{lbf-ft}{lbm°R})\times 574.67°R"

Unit of bath side is balanced

Diametr D=2.537 ft

Height "H=3\times D=3\times 2.537"

H=7.6115ft